body of mass 0.25 KG moving with velocity 12 m/s is stopped by applying a force of 0.6 N calculate the time taken to stop the body. Also calculate the impulse of this force.


1
Expert's answer
2021-08-05T14:17:43-0400

"p=mv=3~\frac {kg\cdot m}s,"

"F=\frac pt,\implies t=\frac pF=5~s."